Use the quadratic formula:
3x^2 - 2x + 1 = 0

Answer by checkley71(8403)

X=(-B+-SQRT[B^2-4AC])/2A
X=(2+-SQRT2*2-4*3*1])/2*3
X=(2+-SQRT4-12])/6
X=(2+-SQRT[-8])/6
X=(2+-2.8284i)/6
X=2/6+2.8284i/6
X=1/3+.47i SOLUTION
X=1/3-.47i SOLUTION